OpenStudy (anonymous):

Find the area of three-fifths of a circle with radius 15 ft. Enter the numerical answer only, rounded to the nearest tenth.

OpenStudy (anonymous):

Find the area of a circle with a radius of 15 ft and multiply it by $\frac{3}{5}$

OpenStudy (kainui):

So 3/5 of the area of a circle, you must know that the area of a circle is A=pi*r^2.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what would r be?

OpenStudy (anonymous):

would the answer be 424.115

OpenStudy (anonymous):

\[\frac{3}{5}\pi 15^2=135 \pi =424.1 \text{ ft}^2 \]

OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm? Enter the numerical answer only. would this be 11.5??

OpenStudy (chris911):

@robtobey nobody gave u an medal lol

OpenStudy (anonymous):

Thank you for the medals.